OpenStudy (anonymous):

Hi, I am new to this website :) I'm currently taking an online trig class on De Moivre's theorem and I don't understand it at all! The question is: Write each expression in the standard form for a complex number, a+bi. 1. [3cos(27))+isin(27)]^5 2. [2(cos(40))+isin(40)]^6 any info regarding this question would be extremely helpful! Thanks!

jimthompson5910 (jim_thompson5910):

De Moivre's theorem says that if you have z = r*[ cos(theta) + i*sin(theta) ] then z^n = r^n*[ cos(n*theta) + i*sin(n*theta) ]

jimthompson5910 (jim_thompson5910):

so that means [3(cos(27)+isin(27)]^5 3^5*[cos(5*27)+isin(5*27)] 243*[cos(135)+isin(135)] I'll let you finish up. Use the unit circle from here and simplify

OpenStudy (anonymous):

Oh!! That makes so much more sense now, thank you! :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

I have one more question regarding the same there, how do I find the complex cube roots of a function?

OpenStudy (anonymous):

Same *theorem*

jimthompson5910 (jim_thompson5910):

what's the question

OpenStudy (anonymous):

Find the complex cube roots of 8(cos(4pi/5) + isin(4pi/5))

jimthompson5910 (jim_thompson5910):

well one cube root (there are 3 total) is found by taking the cube root of 8 to get 2 then you divide each argument by 3, so one cube root is 2*[cos(4pi/15) + i*sin(4pi/15)]

jimthompson5910 (jim_thompson5910):

actually, i found a better formula

jimthompson5910 (jim_thompson5910):

if z = r*[ cos(theta) + i*sin(theta) ] then z^(1/n) = r^(1/n)*[ cos( (theta+360*k)/n ) + i*sin( (theta+360*k)/n ) ] the notation z^(1/n) means the nth root of z the term k is a constant that ranges from 0 to n

jimthompson5910 (jim_thompson5910):

so for instance, in this case r = 8 theta = 4pi/5 n = 3 then you'll go through 3 values of k: k = 0, k = 1, k = 2 sorry I meant to say that k ranges from 0 to n-1

jimthompson5910 (jim_thompson5910):

oh and replace 360 with 2pi if you're using radians

OpenStudy (anonymous):

Ok so the one part I don't get is how you got 3 exactly? Not to sound dumb lol I understood everything else you said

OpenStudy (anonymous):

Oh it is because n is the root right! Sorry

jimthompson5910 (jim_thompson5910):

yes exactly, cube roots will have n = 3 square roots will use n = 2 etc etc

OpenStudy (anonymous):

Right, so I will have 3 different solutions all together right? For each root from 0-3?

jimthompson5910 (jim_thompson5910):

0-2 actually, i made that mistake myself lol

jimthompson5910 (jim_thompson5910):

so the first cube root will have you use r = 8 theta = 4pi/5 n = 3 k = 0 the second cube root will use the same values but k will be different r = 8 theta = 4pi/5 n = 3 k = 1 the last cube root will only change k as well r = 8 theta = 4pi/5 n = 3 k = 2

OpenStudy (anonymous):

And just to clarify that I am solving these correctly, should my answer be just one number or a function or something? Sorry for all these questions lol

jimthompson5910 (jim_thompson5910):

its ok, questions are good

jimthompson5910 (jim_thompson5910):

with cube roots, you'll have 3 answers total

jimthompson5910 (jim_thompson5910):

it's analogous to saying that square roots produce two answers

jimthompson5910 (jim_thompson5910):

ex: if x^2 = 25 then x = -5 or x = 5

jimthompson5910 (jim_thompson5910):

turns out that if x^3 = 216, then x = 6, x = -3-3i*sqrt(3), x = -3+3i*sqrt(3)

jimthompson5910 (jim_thompson5910):

this is because the fundamental theorem of algebra states that if f(x) is a polynomial function of degree n, then f(x) will have n complex roots